# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v2

@@ 3.3.1 Traditional Chebyshev filtering approach with inexact matrix-products involving $\textbf{D}^{-1}$
$\underline{\textbf{Y}}_{k+1}^{(i)}=a_{k}{\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}+b_{k}\underline{\textbf{Y}}_{k}^{(i)}+c_{k}\underline{\textbf{Y}}_{k-1}^{(i)}$ (38)
with the initial conditions $\underline{\textbf{Y}}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\underline{\textbf{Y}}_{1}^{(i)}=\sigma_{1}\frac{\textbf{D}^{-1}\textbf{A}-c\textbf{I}}{e}\textbf{X}^{(i)}$. This allows us to write a recurrence relation for $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
$\bm{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\bm{\Delta}_{k}^{(i)}+b_{k}\bm{\Delta}_{k}^{(i)}+c_{k}\bm{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{B}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}\right)$ (39)
with the initial conditions $\bm{\Delta}_{0}^{(i)}=\bm{\Delta}_{1}^{(i)}=0$.
@@
The spectral norm of the error $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using inexact matrix products of the recurrence relation Eq. 3 involving $\textbf{D}^{-1}$ satisfies $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
Proof.
We argue by induction that $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ where $\eta_{k}$ is some finite constant that depends on $k$. To this end, assuming the induction hypothesis for $k=r-1$ and $k=r$, we have $\lVert\bm{\Delta}_{r-1}^{(i)}\rVert\,\leq\zeta\eta_{r-1}$ and $\lVert\bm{\Delta}_{r}^{(i)}\rVert\,\leq\zeta\eta_{r}$. Subsequently, the following inequality is true for $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ from Eq. 39
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ $\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}\zeta+\left\lvert b_{r}\right\rvert\,\eta_{r}\zeta+\lvert c_{r}\rvert\,\eta_{r-1}\zeta+\left\lvert a_{r}\right\rvert\,\lVert{\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{r}^{(i)}}-\textbf{B}^{-1}\textbf{A}\underline{\textbf{Y}}_{r}^{(i)}\rVert\,$ (40)
$\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}\zeta+\left\lvert b_{r}\right\rvert\,\eta_{r}\zeta+\lvert c_{r}\rvert\,\eta_{r-1}\zeta+\left\lvert a_{r}\right\rvert\,\lVert\textbf{A}\rVert\,\lVert\underline{\textbf{Y}}_{r}^{(i)}\rVert\,\zeta$ (41)
We further note that $\lVert\underline{\textbf{Y}}_{r}^{(i)}\rVert\,=\lVert\bm{\Delta}_{r}^{(i)}+\textbf{Y}_{r}^{(i)}\rVert\,\leq\zeta\eta_{r}+\lVert\textbf{Y}_{r}^{(i)}\rVert\,\leq\zeta\eta_{r}+\lVert C_{r}(\textbf{H})\textbf{X}^{(i)}\rVert\,\leq\zeta\eta_{r}+\lVert C_{r}(\textbf{H})\rVert\,$, and consequently we can write
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ $\leq\left[\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}+\left\lvert b_{r}\right\rvert\,\eta_{r}+\lvert c_{r}\rvert\,\eta_{r-1}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,(\zeta\eta_{r}+\lVert C_{r}(\textbf{H})\rVert\,)\right]\zeta$ (42)
Defining $\eta_{r+1}=\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\eta_{r}+\left\lvert b_{r}\right\rvert\,\eta_{r}+\lvert c_{r}\rvert\,\eta_{r-1}+\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,(\zeta\eta_{r}+\lVert C_{r}(\textbf{H})\textbf{B}^{-\frac{1}{2}}\rVert\,)$ we have $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,\leq\zeta\eta_{r+1}$from the above equation. Hence using induction, we argue that $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ for $k=0,1,\dots,p$
@@
We now discuss this result in the context of Eq. 6. Without loss of generality, we assume that our initial guess at $i=0$ satisfies Eq. 6. While Eq. 6 is satisfied this allows for $\angle(\mathcal{S}^{(i)},\mathcal{S})$ to decrease as $i$ increases, but we note that the RHS of Eq. 6 does not monotonically decrease with $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and in fact increases as $\angle(\mathcal{S}^{(i)},\mathcal{S})$ approaches 0. Thus, it stands to reason that beyond a certain value of $\angle(\mathcal{S}^{(i)},\mathcal{S})$ Eq. 6 no longer holds, and we cannot approach the wanted eigenspace beyond this point. Assuming that $\zeta\eta_{p}<<\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ we estimate that the closest angle that can be achieved is
@@
$\angle(\mathcal{S}^{(i)},\mathcal{S})\approx\frac{\zeta\eta_{p}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}$ (43)
@@ 3.3.2 Proposed residual-based Chebyshev filtering approach for generalized eigenproblem
We now propose a residual-based reformulation of the recurrence relation defined in Eq. 3. To this end we redefine the residual $\textbf{R}_{k}^{(i)}=\textbf{B}^{-1}\textbf{D}(C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\bm{\Lambda}^{(i)}))=\textbf{B}^{-1}\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)})$ for $k=0,\dots,p$, where we have defined $\bm{\Lambda}_{k}^{(i)}=C_{k}(\bm{\Lambda}^{(i)})$.
@@
The recurrence relation given by Eq. 3 can be reformulated in terms of the $\textbf{B}\textbf{R}_{k}^{(i)}=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)})$ as
$\textbf{B}\textbf{R}_{k+1}^{(i)}=a_{k}\textbf{D}\textbf{H}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}+b_{k}\textbf{B}\textbf{R}_{k}^{(i)}+c_{k}\textbf{B}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{D}\textbf{B}^{-1}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$ (44)
where $\textbf{B}\textbf{R}_{0}^{(i)}=0$ and $\textbf{B}\textbf{R}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{D}\textbf{B}^{-1}\textbf{R}^{(i)}$ with $\textbf{R}^{(i)}=\textbf{A}\textbf{X}^{(i)}-\textbf{B}\textbf{X}^{(i)}\bm{\Lambda}^{(i)}$.
Proof.
The recurrence relation for $\bm{\Lambda}_{k}^{(i)}=C_{k}(\bm{\Lambda}^{(i)})$ can be written as
$\bm{\Lambda}_{k+1}^{(i)}=a_{k}\bm{\Lambda}_{k}^{(i)}\bm{\Lambda}^{(i)}+b_{k}\bm{\Lambda}_{k}^{(i)}+c_{k}\bm{\Lambda}_{k-1}^{(i)}$
@@
$\textbf{Y}_{k+1}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k+1}^{(i)}$ $=a_{k}(\textbf{H}\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)}\bm{\Lambda}^{(i)})+b_{k}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)})+c_{k}(\textbf{Y}_{k-1}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)})$
$\implies\textbf{D}^{-1}\textbf{B}\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{H}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}+b_{k}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}+c_{k}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{B}^{-1}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$
$\implies\textbf{B}\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{D}\textbf{H}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}+b_{k}\textbf{B}\textbf{R}_{k}^{(i)}+c_{k}\textbf{B}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{D}\textbf{B}^{-1}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$
After computing $\textbf{B}\textbf{R}_{p}^{(i)}$ using this recurrence relation we can now evaluate $\textbf{Y}_{p}^{(i)}$ using the relation $\textbf{Y}_{p}^{(i)}=\textbf{D}^{-1}\textbf{B}\textbf{R}_{p}^{(i)}+\textbf{X}^{(i)}\bm{\Lambda}_{p}^{(i)}$. We note that using approximation $\textbf{D}\textbf{B}^{-1}\approx\textbf{I}$ we can write the following recurrence relation
$\textbf{B}\underline{\textbf{R}}_{k+1}^{(i)}=a_{k}{\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}}+b_{k}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}+c_{k}\textbf{B}\underline{\textbf{R}}_{k-1}^{(i)}+a_{k}\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}$ (45)
with $\textbf{B}\textbf{R}_{0}^{(i)}=0$ and $\textbf{B}\textbf{R}_{1}^{(i)}=\frac{2\sigma_{2}}{e}\textbf{R}^{(i)}$ and consequently, we have $\underline{\textbf{Y}}_{k}^{(i)}=\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}+\textbf{X}^{(i)}\bm{\Lambda}_{k}^{(i)}$. Note that this recurrence relation does not require the evaluation $\textbf{B}^{-1}$. We can now write a recurrence relation for $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}=\textbf{D}^{-1}\textbf{B}(\underline{\textbf{R}}_{k}^{(i)}-\textbf{R}_{k}^{(i)})$ as
$\bm{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\bm{\Delta}_{k}^{(i)}+b_{k}\bm{\Delta}_{k}^{(i)}+c_{k}\bm{\Delta}_{k-1}^{(i)}+a_{k}\left(\textbf{I}-\textbf{D}\textbf{B}^{-1}\right)\left(\textbf{R}^{(i)}\bm{\Lambda}_{k}^{(i)}+\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}\right)$ (46)
where $\bm{\Delta}_{0}^{(i)}=0$ and $\bm{\Delta}_{1}^{(i)}=\frac{2\sigma_{2}}{e}(\textbf{I}-\textbf{D}\textbf{B}^{-1})\textbf{R}^{(i)}$.
@@
The spectral norm of the error $\bm{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using recurrence relation given by Eq. 45 satisfies $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$ where $\tilde{\eta}_{k}$ is some finite constant that depends on $k$ and $\tilde{\zeta}=\lVert\textbf{I}-\textbf{D}\textbf{B}^{-1}\rVert\,$.
Proof.
We argue by induction that $\lVert\bm{\Delta}_{k}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ where $\tilde{\eta}_{k}$ is some non-zero finite constant that depends on $k$. To this end, assuming the induction hypothesis for $k=r-1$ and $k=r$, we have $\lVert\bm{\Delta}_{r}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{r}\lVert\textbf{R}^{(i)}\rVert\,$ and $\lVert\bm{\Delta}_{r-1}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{r-1}\lVert\textbf{R}^{(i)}\rVert\,$. Subsequently, the following inequality is true for $\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ from Eq. 46
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,$ $\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert a_{r}\right\rvert\,\lVert\left(\textbf{I}-\textbf{D}\textbf{B}^{-1}\right)\left(\textbf{R}^{(i)}\bm{\Lambda}_{r}^{(i)}+\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{r}^{(i)}\right)\rVert\,$

### v3

@@ 3.3.1 Traditional Chebyshev filtering approach with inexact matrix-products involving $\textbf{D}^{-1}$
(21) $\underline{\textbf{Y}}_{k+1}^{(i)}=a_{k}{\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}+b_{k}\underline{\textbf{Y}}_{k}^{(i)}+c_{k}\underline{\textbf{Y}}_{k-1}^{(i)}$
with the initial conditions $\underline{\textbf{Y}}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\underline{\textbf{Y}}_{1}^{(i)}=\sigma_{1}\frac{\textbf{D}^{-1}\textbf{A}-c\textbf{I}}{e}\textbf{X}^{(i)}$.
@@
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using inexact matrix products of the recurrence relation Eq. 2 involving $\textbf{D}^{-1}$ satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ is some finite constant that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
Proof 3.20.
Using Eq. 2 and Eq. 21 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
(22) $\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\left({\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{B}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}\right)$
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=0$ and $\boldsymbol{\Delta}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{X}^{(i)}$. We note that this recurrence relation is of the same form as that of Eq. 8 in Lemma 3.5 with $\textbf{e}_{k}^{(i)}={\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{B}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}$. We now have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{D}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{B}^{-1}\textbf{A}\underline{\textbf{Y}}_{k}^{(i)}\rVert\,\leq\zeta\lVert\textbf{A}\rVert\,\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,$. We further note that
$\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\textbf{X}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\zeta\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\zeta\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$. Using Lemma 3.5 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{X}^{(i)}\rVert\,\leq\lvert\frac{\sigma_{1}}{e}\rvert\,\zeta$, $h_{0}=\zeta\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=\zeta\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\eta_{k}$ with $\eta_{k}=g_{k-1}^{0}\lvert\frac{\sigma_{1}}{e}\rvert\,+\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}$.
@@
We now discuss this result in the context of Eq. 5. Without loss of generality, we assume that our initial guess at $i=0$ satisfies Eq. 5 and this allows for $\angle(\mathcal{S}^{(i)},\mathcal{S})$ to decrease as $i$ increases. However, we note that the RHS of Eq. 5 does not monotonically decrease with $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and in fact increases as $\angle(\mathcal{S}^{(i)},\mathcal{S})$ approaches 0. Thus, it stands to reason that beyond a certain value of $\angle(\mathcal{S}^{(i)},\mathcal{S})$ Eq. 5 no longer holds, and we cannot approach the wanted eigenspace beyond this point. Assuming that $\zeta\eta_{p}<<\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ we estimate that the closest angle that can be achieved is
@@
(23) $\angle(\mathcal{S}^{(i)},\mathcal{S})\approx\frac{\zeta\eta_{p}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}$
@@ 3.3.2 Proposed residual-based Chebyshev filtering approach for generalized eigenproblem
We now propose a residual-based reformulation of the recurrence relation defined in Eq. 2. To this end we redefine the residual $\textbf{R}_{k}^{(i)}=\textbf{B}^{-1}\textbf{D}(C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)}))=\textbf{B}^{-1}\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})$ for $k=0,\dots,p$, where we have defined $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$.
@@
The recurrence relation given by Eq. 2 can be reformulated in terms of the $\textbf{B}\textbf{R}_{k}^{(i)}=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})$ as
(24) $\textbf{B}\textbf{R}_{k+1}^{(i)}=a_{k}\textbf{D}\textbf{H}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}+b_{k}\textbf{B}\textbf{R}_{k}^{(i)}+c_{k}\textbf{B}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{D}\textbf{B}^{-1}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
where $\textbf{B}\textbf{R}_{0}^{(i)}=0$ and $\textbf{B}\textbf{R}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{D}\textbf{R}^{(i)}$ with $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$.
Proof 3.22.
The recurrence relation for $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ can be written as
$\boldsymbol{\Lambda}_{k+1}^{(i)}=a_{k}\boldsymbol{\Lambda}_{k}^{(i)}\boldsymbol{\Lambda}^{(i)}+b_{k}\boldsymbol{\Lambda}_{k}^{(i)}+c_{k}\boldsymbol{\Lambda}_{k-1}^{(i)}$
@@
$\textbf{B}\textbf{R}_{k+1}^{(i)}$ $=a_{k}\textbf{D}\textbf{H}\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}+b_{k}\textbf{B}\textbf{R}_{k}^{(i)}+c_{k}\textbf{B}\textbf{R}_{k-1}^{(i)}+a_{k}\textbf{D}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
After computing $\textbf{B}\textbf{R}_{p}^{(i)}$ using this recurrence relation we can now evaluate $\textbf{Y}_{p}^{(i)}$ using the relation $\textbf{Y}_{p}^{(i)}=\textbf{D}^{-1}\textbf{B}\textbf{R}_{p}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{p}^{(i)}$. We note that using approximation $\textbf{D}\textbf{B}^{-1}\approx\textbf{I}$ we can write the following recurrence relation
(25) $\textbf{B}\underline{\textbf{R}}_{k+1}^{(i)}=a_{k}{\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}}+b_{k}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}+c_{k}\textbf{B}\underline{\textbf{R}}_{k-1}^{(i)}+a_{k}\textbf{B}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
with $\textbf{B}\textbf{R}_{0}^{(i)}=0$ and $\textbf{B}\textbf{R}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{B}\textbf{R}^{(i)}$ and consequently, we have $\underline{\textbf{Y}}_{k}^{(i)}=\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$. Note that this recurrence relation does not require the evaluation of $\textbf{B}^{-1}$.
@@
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using recurrence relation given by Eq. 25 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$ where $\tilde{\eta}_{k}$ is some finite constant that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$.
Proof 3.24.
Using Eq. 24 and Eq. 25 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}=\textbf{D}^{-1}\textbf{B}(\underline{\textbf{R}}_{k}^{(i)}-\textbf{R}_{k}^{(i)})$ as
$\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}\!+\!a_{k}\left(\textbf{D}^{-1}-\textbf{B}^{-1}\right)\left(\textbf{B}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}+\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}\right)$
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=0$ and $\boldsymbol{\Delta}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{B}\textbf{R}^{(i)}$. We note that this recurrence relation is of the same form as that of Eq. 8 in Lemma 3.5 with
$\textbf{e}_{k}^{(i)}=\left(\textbf{D}^{-1}-\textbf{B}^{-1}\right)\left(\textbf{B}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}+\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}\right)$